Q. a gun of mass "M" fires a bullet of mass "M/n".if the energy of the bullet is "E",then the energy spent in firing the bullet is ?


ans: (1+1/n)E


please help!!!

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

Let p be momentum of bullet; p2 =2(M/n)E, therefore p=√(2(M/n)E). from law of conservation of momentum, this is equal to momentum og gun(P). therefore, P= √(2Me),(e is energy of gun), p = P implies= √(2(M/n)E) = =√(2Me). therefore e=E/n

therefore total energy = energies of gun+bullet = E+(E/n) =(1+ (1/n))E

'Win exciting gifts by answering the questions on Discussion Forum.